What is virial theorem


1
Expert's answer
2021-01-25T14:03:37-0500

The virial theorem is a relation that connects the average kinetic energy of a system of particles with the forces acting in it. The virial G for the set of N point particles in mechanics is defined as:


"G=\sum{\overrightarrow{p}_k\cdot\overrightarrow{r}_k}" ,


where "\overrightarrow{p}_k" and "\overrightarrow{r}" are the spatial vectors of momentum coordinates and coordinates for the k-th particle. For a stable system bound by potential forces, the virial theorem is


"<T>=-\frac{1}{2}\sum{\overrightarrow{F}_k\cdot\overrightarrow{r}_k}"


where "<T>" is the average total kinetic energy; "\overrightarrow{F}_k" is the force acting on the k-th particle.